# Do assumptions about the central density of subhaloes affect dark matter annihilation and lensing calculations?

### v2

@@ 5 Lensing
where $\kappa$ is the convergence and $\gamma$ is the shear. The corresponding convergence and shear profiles are defined as
@@
for axisymmetric haloes. In the case of multiple lenses, the total amplification matrix can then be calculated by summing the contributions of each lens.
The projected density profile $\Sigma(R)$ can be calculated from the density profile,
@@
and the mean projected surface mass
@@
with the projected mass $M_{\rm proj}(R)$ is given by:
@@
In Figure 12, the projected density profiles (top) show similar trends to the 3D density profiles, with $\Sigma(r>r_{t})=0$ (except for H03). The shear profiles (bottom) are approximately constant within the tidal radius and then follow $\Delta\Sigma\propto R^{-2}$. The discontinuity in the slope of the ST model is evident in the shear profiles; the shear is finite but not differentiable at the cutoff radius, as discussed in [3]. Figure 13 shows the evolution of $\Sigma$ and $\Delta\Sigma$, calculated within $0.5\,r_{\rm unit}$ and $1\,r_{\rm unit}$, as mass is lost. Initially, the EinHighT lens models have slightly higher projected densities than the other models, but generally, all models agree to within 10–20 per cent, except at the discontinuity where the tidal radius is close to the aperture size in the ST model.
@@ 5.1 Model Comparisons
To understand better how the convergence and shear depend on the assumed subhalo profile and stripping model, in Figure 14, we compare radial profiles and residuals for the different models. Between 0.1 $r_{-2}$ and $r_{t}$, EinHighT and EinLowT $\Sigma$ differ from the NFWT profile by up to 20 per cent, while the shear profiles differ from NFWT profiles by up to 50 per cent. The difference is largest at very small radii, with the EinLowT profile having higher convergence and sheer values, due to the increased density in the centre of the profile. We also compare the energy-truncated NFW profiles to the commonly used H03 and ST approximations. The H03 model underestimates the convergence and shear profiles at small radii, though differences between H03 and NFWT are comparable to differences found between the different initial profiles, assuming energy truncation. The ST model closely matches the NFWT model at small radii but is less extended.
@@
Figure 14: The projected density profile, $\Sigma$, (top) and $\Delta\Sigma\equiv\bar{\Sigma}-\Sigma$ (third row) for the Energy Truncation model applied to three different profiles (NFWT, EinLowT and EinLowT), the stripping model from hayashi2003 [39] applied to an NFW profile (H03; dotted line), and a sharply truncated NFW profile (ST; dashed line). Each column corresponds to a different bound mass, as indicated. The second and last rows show the ratio of $\Sigma$ and $\Delta\Sigma$, respectively, to the NFWT model. The grey box indicates radii that are typically unresolved in isolated simulations. There are significant differences between the three initial halo models, but the ST approximation does a good job of approximating the NFWT profiles, except for $\Delta\Sigma$ measured close to the truncation radius.
@@
Overall, this suggests that errors in the convergence and shear profiles are dominated by assumptions of the initial profile rather than the tidal mass-loss model; i.e., while an ST model does a fairly good job of predicting the convergence and shear of subhalo profiles, different values of the Einasto $\alpha$ value can introduce larger differences. Thus, we suspect that using updated tidal models (such as the Energy Truncation model) is unlikely to improve models of satellite lensing.
@@
Previous work typically finds that different lens models introduce an error of $\sim 10$ per cent in mass estimates [e.g. 57, 3, 73]. Additionally, it has been found that even strong tidal truncation has a negligible effect on fitted subhalo parameters [$M_{200}$ and concentration; 59] — however, $M_{200}$ is not necessarily a good estimation of the bound subhalo mass. We explore the effect of profile assumptions on subhalo mass estimates in Figure 15. We assume that all projected density profiles can be well-described by a sharp truncation of the NFW model (i.e. the ST model), and fit the free parameters $r_{s}$, $\rho_{0}$, and $r_{t}$. We then calculated the total mass by integrating the projected density profile. We find that the relative error in the subhalo mass typically ranges from 5–10 per cent.
@@
Figure 15: Top: The projected density profile, $\Sigma$ for the Energy Truncation model applied to three different profiles (NFWT, EinLowT, and EinLowT), the stripping model from hayashi2003 [39] applied to an NFW profile (H03; dotted line), and a sharply truncated NFW profile (ST; dashed line). Each column corresponds to a different bound mass, as indicated. The thick transparent lines are the best-fit ST models. We also show the percent error (PE) in the mass estimate calculated as $(M_{\rm ST}-M_{\rm true})/M_{\rm true}\times 100\%$. Bottom: residuals in $\Sigma$ compared to the best-fit ST model.
@@ 5.2 Model Sensitivity in Lensing: Summary
We find that while there is a significant (30% or more) difference in the projected density and shear between NFW and Einasto profiles with different values of $\alpha$, the tidal stripping model is less important. The central values of lensing observables predicted by the Energy Truncation model are within $\sim$10of those predicted by the ST (sharp truncation) approximation. In particular, subhalo mass estimates based on ST are accurate to within 10 per cent, regardless of the mass profile assumed. Thus, for lensing applications, modelling results are not sensitive to assumptions about tidal stripping models.
@@ 6 Discussion
In this work, we have considered the implications of the uncertainty in subhalo density profiles and tidal evolution for calculations of the dark matter annihilation rate and lensing signals. Overall, we find that the annihilation rate and boost factor are very sensitive to both the subhalo profile and mass loss modelling; at small bound mass ratios, different assumptions lead to annihilation signals that differ by an order of magnitude. While there were also varying predictions in the quantities measured in subhalo lensing, we find that these are generally much less sensitive.
@@
As discussed in [26, 27], since subhalo evolution is primarily understood from idealized simulations (due to resolution constraints), predictions for subhalo evolution depend sensitively on the initial halo models assumed in simulations. Here, we have considered Einasto profiles spanning the range of $\alpha$ values one might expect to exist, and also NFW profiles, which are useful for comparison with earlier work. Since lower Einasto $\alpha$ values correspond to steeper density profiles, the results in this paper from the EinLow model are the most relevant for the case of microhaloes, the smallest structures to form in the early universe, whose size is determined by the free-streaming scale of the dark matter particles (although the peak-height dependence of the Einasto shape parameter complicates this statement slightly).
@@
Assuming dark matter particles have a mass of 100 GeV, the mass of the smallest microhaloes is approximately Earth mass [e.g. 22]. It is expected that the central density profile of these microhaloes is steeper than that found in larger haloes and thus may contribute greatly to the dark matter annihilation signal [e.g. 79]. We found the concentration of EinLowT profiles increases rapidly during tidal stripping and has larger annihilation signals than NFWT profiles at a fixed mass. Interestingly, in [29], it was found that Einasto profiles with low $\alpha$ values were the only ones that decreased in concentration after major mergers. These results all suggest that profiles with cuspier centres (as found in the earliest haloes) evolve differently.
@@
[45] found that microhaloes may have much higher densities than expected from extrapolations of the low-redshift concentration–mass–redshift relations, and [68] showed that if these densities are conserved, concentrations of the smallest haloes are still uncertain by a factor of $\sim 5$ when extrapolated to low redshift (considering the redshift evolution of the virial radius). [68] show that this uncertainty in concentration translates to an increased boost factor by up to two orders of magnitude. In our work, we find that the Energy Truncation model results in boost factors that are approximately 5 times higher at a given effective concentration (Figure 9). These results offer separate evidence that the boost factors of the smallest haloes may be greatly underestimated.
@@
While we have focused on dark matter annihilation and lensing applications, subhaloes can also be used to study the nature of dark matter through stellar streams created by disrupted dwarf galaxies. For example, stream morphology can help constrain the potential of the host dark matter halo [e.g. 12], and features in streams caused by perturbations from other subhaloes can be used to place constraints on dark matter properties [e.g. 8]. In particular, there has been a lot of interest in studying the GD–1 stellar stream [6], which may have been perturbed by a compact dark matter subhalo. Other applications include semi-analytic models of galaxy formation, of which subhalo evolution models are often a key component [e.g. 90, 47]. We leave the implications of conserved central density on these applications to future work.
@@
A major assumption in this work is that the Energy Truncation model is an accurate representation of subhalo evolution. However, this model has several assumptions that are important to address. First, the infalling subhaloes are assumed to be spherical and isotropic. We expect that as the material is stripped off of subhaloes, the subhalo quickly becomes spherical and isotropic [26], and therefore this assumption is approximately valid. However, we plan to explore tidal stripping on non-spherical systems in future work. One way to create non-spherical initial conditions for idealized simulations is by using the output of major merger simulations, as proposed in [60]; in future work, we plan to use the simulations presented in [28, 29] to explore stripping of non-spherical systems.
@@
Perhaps the most important assumption made in the Energy Truncation model is that subhaloes only evolve due to tidal fields, and there is no interaction with a baryonic component; extensive work has shown that baryonic matter influences dwarf galaxy properties [as reviewed in 7]. It has been proposed that dwarf galaxy haloes can transform from cusps to cores based on the stellar-to-halo mass ratio, $M_{*}/M_{200}$ [70], where the cusp is protected from supernova feedback if $M_{*}/M_{200}\lesssim 5\times 10^{-4}$ [21], which has been supported by several measurements [72, 13]. These considerations may be relevant for weak lensing studies of substructure since they centre on subhaloes with associated stellar mass by definition. For annihilation studies, on mass scales of $M_{200}<10^{10}\,M_{\odot}$ below the dwarf galaxy range, star formation should be greatly suppressed by photo-heating [e.g. 41, 66]. Therefore, it is expected that dark-matter-dominated systems (down to putative Earth-mass-sized haloes) should not be affected by baryonic feedback. These dark matter-dominated systems may contribute greatly to an annihilation signal.
@@ 7 Conclusion
Our best hope for constraining the properties of dark matter through astrophysical observations comes from tests at the highest dark matter densities. Important examples include searches for a gamma-ray or other signal from dark matter annihilation, and tests of substructure in gravitational lenses. The inferences derived from observations depend strongly on the assumed clustering of dark matter on the smallest scales. Unfortunately, this remains unclear from theory; while there has been progress in understanding the average density profile of isolated haloes, including the smallest ones, it remains unclear how this profile evolves once these become subhaloes. Idealized cases of single mergers have been simulated at increasing resolution, but remain susceptible to numerical effects which can cause increased mass loss [e.g. 83, 82] and decreased central densities [e.g. 33, 27]. To attempt to overcome numerical limitations, we have used a theoretically motivated, universal Energy Truncation method to study subhalo evolution. We quantify the effect of underestimating the central density of tidally-stripped haloes on annihilation signals and lensing profiles of individual subhaloes.
For lensing mass experiments, we conclude that the details of tidal stripping are less important than assumptions about the original density profile. Thus, recent estimates of the uncertainties in mass introduced by different profiles (approximately 10 per cent) are likely reliable. On the other hand, annihilation constraints are extremely sensitive to uncertainties both in subhalo profiles and in the details of tidal stripping. In particular, our Energy Truncation method predicts annihilation signals that are an order of magnitude larger than some previous estimates. Our initial calculations indicate that the total Milky Way annihilation signal may differ by a factor of five or more depending on the subhalo model used; however, the full annihilation calculation is complicated, and beyond the scope of this paper. Future work needs to be done to fully quantify the uncertainties in current dark matter annihilation constraints in light of updated subhalo evolution models.
@@ Acknowledgments
NED acknowledges support from NSERC Canada, through a postgraduate scholarship. JET acknowledges financial support from NSERC Canada, through a Discovery Grant.
Software: numpy [38], matplotlib [42], scipy [84], Colossus [24].

### v3

@@ 5 Lensing
where $\kappa$ is the convergence and $\gamma$ is the shear. In the case of multiple lenses, the total amplification matrix can then be calculated by summing the contributions of each lens.
The radially-averaged convergence and shear profiles are defined as
@@
and the projected density profile $\Sigma(R)$ can be calculated from the density profile,
@@
while the mean projected surface mass density is
@@
given the projected mass:
@@
In Figure 12, the projected density profiles (top) show similar trends to the 3D density profiles, with $\Sigma(r>r_{t})=0$ (except for H03). The shear profiles (bottom) are approximately constant within the tidal radius and then follow $\Delta\Sigma\propto R^{-2}$. The discontinuity in the slope of the ST model is evident in the shear profiles; the shear is finite but not differentiable at the cutoff radius, as discussed in [15]. Figure 13 shows the evolution of $\Sigma$ and $\Delta\Sigma$, calculated within $0.5\,r_{\rm unit}$ and $1\,r_{\rm unit}$, as mass is lost. Initially, the EinCoreT lens models have slightly higher projected densities than the other models, but generally, all models agree to within 10–20 per cent, except at the discontinuity where the tidal radius is close to the aperture size in the ST model.
@@ 5.1 Model comparisons
To understand better how the convergence and shear depend on the assumed subhalo profile and stripping model, in Figure 14, we compare radial profiles and residuals for the different models. Between 0.1 $r_{-2}$ and $r_{t}$, EinCoreT and EinCuspT mean projected density, $\Sigma$, differ from the NFWT profile by up to 20 per cent, while the shear profiles differ from NFWT profiles by up to 50 per cent. The difference is largest at very small radii, with the EinCuspT profile having higher convergence and sheer values, due to the increased density in the centre of the profile. We also compare the energy-truncated NFW profiles to the commonly used H03 and ST approximations. The H03 model underestimates the convergence and shear profiles at small radii, though differences between H03 and NFWT are comparable to differences found between the different initial profiles, assuming energy truncation. The ST model closely matches the NFWT model at small radii but is less extended.
@@
Figure 14: The projected density profile, $\Sigma$, (top) and $\Delta\Sigma\equiv\bar{\Sigma}-\Sigma$ (third row) for the Energy Truncation model applied to three different profiles (NFWT, EinCoreT and EinCuspT), the stripping model from [16] applied to an NFW profile (H03; dotted line), and a sharply truncated NFW profile (ST; dashed line). Each column corresponds to a different bound mass, as indicated. The second and last rows show the ratio of $\Sigma$ and $\Delta\Sigma$, respectively, to the NFWT model. The grey box indicates radii that are typically unresolved in isolated simulations. There are significant differences between the three initial halo models, but the ST model does a good job of approximating the NFWT profiles, except for $\Delta\Sigma$ measured close to the truncation radius.
@@
Overall, this suggests that errors in the convergence and shear profiles are dominated by assumptions of the initial profile rather than the tidal mass-loss model; i.e., while an ST model does a fairly good job of predicting the convergence and shear of subhalo profiles, different values of the Einasto $\alpha$ parameter can introduce larger differences. Thus, we suspect that using updated tidal models (such as the Energy Truncation model) is unlikely to improve satellite lensing calculations significantly.
@@
Previous work typically finds that different lens models introduce an error of $\sim 10$ per cent in mass estimates [e.g. 75, 15, 76]. Additionally, it has been found that even strong tidal truncation has a negligible effect on fitted subhalo parameters [$M_{200}$ and concentration; 77]—however, $M_{200}$ is not necessarily a good estimation of the bound subhalo mass. We explore the effect of profile assumptions on subhalo mass estimates in Figure 15. We assume that all projected density profiles can be well-described by a sharp truncation of the NFW model (i.e. the ST model), and fit the free parameters $r_{s}$, $\rho_{0}$, and $r_{t}$. We then calculated the total mass by integrating the projected density profile. We find that the relative error in the subhalo mass typically ranges from 5–10 per cent.
@@
Figure 15: Top: The projected density profile, $\Sigma$ for the Energy Truncation model applied to three different profiles (NFWT, EinCoreT, and EinCuspT), the stripping model from [16] applied to an NFW profile (H03; dotted line), and a sharply truncated NFW profile (ST; dashed line). Each column corresponds to a different bound mass, as indicated. The thick transparent lines are the best-fit ST models. We also show the percent error (PE) in the mass estimate calculated as $(M_{\rm ST}-M_{\rm true})/M_{\rm true}\times 100\%$. Bottom: residuals in $\Sigma$ compared to the best-fit ST model.
@@ 5.2 Model sensitivity in lensing: Summary
We find that while there is a significant (30% or more) difference in the projected density and shear between NFW and Einasto profiles with different values of $\alpha$, the tidal stripping model is less important. The central values of lensing observables predicted by the Energy Truncation model are within $\sim$10of those predicted by the ST (sharp truncation) approximation. In particular, subhalo mass estimates based on ST agree to within 10 per cent, regardless of the mass profile assumed. Thus, we conclude that for lensing applications, modelling results are not sensitive to assumptions about tidal stripping models.
@@ 6 Discussion
In this work, we have considered the implications of the uncertainty in subhalo density profiles and tidal evolution for calculations of the dark matter annihilation rate and lensing signals. Specifically, we use the Energy Truncation Model to examine the scenario where cuspy centres are preserved under different subhalo profile assumptions. Overall, we find that the annihilation rate and flux multiplier are very sensitive to both the subhalo profile and mass loss modelling; at small bound mass ratios, different assumptions lead to annihilation signals that differ by an order of magnitude. While the quantities measured in subhalo lensing also vary slightly, we find that these are generally much less sensitive to assumptions about tidal stripping.
@@
As discussed in [44, 25], since subhalo evolution is primarily understood from idealized simulations (due to resolution constraints), predictions for subhalo evolution depend sensitively on the initial halo models assumed in simulations. Here, we have considered Einasto profiles spanning the range of $\alpha$ values one might expect to exist, and also NFW profiles, which are useful for comparison with earlier work.
Since lower Einasto $\alpha$ values correspond to steeper density profiles, the results in this paper from the EinCusp model are the most relevant for the case of microhaloes, the smallest structures to form in the early universe, whose size is determined by the free-streaming scale of the dark matter particles (although the peak-height dependence of the Einasto shape parameter complicates this statement slightly).
@@
Assuming dark matter particles have a mass of 100 GeV, the mass of the smallest microhaloes is approximately Earth mass [e.g. 78]. It is expected that the central density profile of these microhaloes is steeper than that found in larger haloes and thus may contribute greatly to the dark matter annihilation signal [e.g. 33]. We found the concentration of EinCuspT profiles increases rapidly during tidal stripping and has larger annihilation signals than NFWT profiles at a fixed mass. Interestingly, in [51], it was found that Einasto profiles with low $\alpha$ values were the only ones that decreased in concentration after major mergers. These results all suggest that profiles with cuspier centres (as found in the earliest haloes) evolve differently.
@@
For instance, haloes may have much higher densities than expected from extrapolations of the low-redshift concentration–mass–redshift relations [26], and if these densities are conserved, concentrations of the smallest haloes are still uncertain by a factor of $\sim 5$ when extrapolated to low redshift (considering the redshift evolution of the virial radius)[2]. This uncertainty in concentration translates to an increased flux multiplier by up to two orders of magnitude. In our work, we find that the Energy Truncation model results in flux multipliers that are approximately 5 times higher at a given effective concentration (Figure 9). These results offer separate evidence that the flux multipliers of the smallest haloes may be greatly underestimated.
@@
While we have focused on dark matter annihilation and lensing applications, subhaloes can also be used to study the nature of dark matter through stellar streams created by disrupted dwarf galaxies. For example, stream morphology can help constrain the potential of the host dark matter halo [e.g. 79], and features in streams caused by perturbations from other subhaloes can be used to place constraints on dark matter properties [e.g. 80]. In particular, there has been a lot of interest in studying the GD–1 stellar stream [81], which may have been perturbed by a compact dark matter subhalo. Other applications include semi-analytic models of galaxy formation, of which subhalo evolution models are often a key component [e.g. 82, 83]. We leave the implications of conserved central density on these applications to future work.
@@
A major assumption in this work is that the Energy Truncation model is an accurate representation of subhalo evolution in the limit that central density is preserved. However, this model has several assumptions that are important to address. First, the infalling subhaloes are assumed to be spherical and isotropic. We expect that as the material is stripped off of subhaloes, the subhalo quickly becomes spherical and isotropic [44], and therefore this assumption is approximately valid. However, we plan to explore tidal stripping on non-spherical systems in future work. One way to create non-spherical initial conditions for idealized simulations is by using the output of major merger simulations, as proposed in [84]; in future work, we plan to use the simulations presented in [85, 51] to explore stripping of non-spherical systems.
Additionally, the concept of "energy" used in the Energy Truncation model is not uniquely defined in a time-dependent, extended system. The binding energy of a particle depends on the chosen frame of reference, the definition of the potential, and the time at which the calculation is performed. While our energy-based ordering is well defined under the assumptions of spherical symmetry and a slowly evolving potential, it is not strictly conserved in more complex or rapidly varying systems. Nonetheless, numerical experiments show that this approach reliably captures the selective stripping of weakly bound particles under gradual mass loss, lending empirical support to its use in the Energy Truncation framework [46, 43, 44]. However, in scenarios involving rapid or impulsive stripping—such as disk shocking or strong subhalo–disk encounters—energy injection can be non-adiabatic, and the assumptions underlying the Energy Truncation model may break down. In such cases, the model may not accurately describe the evolution of subhalo structure, though it may still serve as a useful intermediate baseline between empirical fits like the H03 model and extreme cases such as sharp truncation.
@@
Perhaps the most important assumption made in the Energy Truncation model is that subhaloes only evolve due to tidal fields, and there is no interaction with a baryonic component; extensive work has shown that baryonic matter influences dwarf galaxy properties [as reviewed in 1]. It has been proposed that dwarf galaxy haloes can transform from cusps to cores based on the stellar-to-halo mass ratio, $M_{*}/M_{200}$ [86], where the cusp is protected from supernova feedback if $M_{*}/M_{200}\lesssim 5\times 10^{-4}$ [87], which has been supported by several measurements [88, 89]. These considerations may be relevant for weak lensing studies of substructure since they centre on subhaloes with associated stellar mass by definition. For annihilation studies, on mass scales below the dwarf galaxy range ( $M_{200}<10^{10}\,M_{\odot}$), star formation should be greatly suppressed by photo-heating [e.g. 90, 91]. Therefore, it is expected that dark-matter-dominated systems (down to putative Earth-mass-sized haloes) should not be affected by baryonic feedback. These dark matter-dominated systems may contribute greatly to an annihilation signal.
@@ 7 Conclusion
Our best hope for constraining the properties of dark matter through astrophysical observations comes from tests at the highest dark matter densities. Important examples include searches for a gamma-ray or other signal from dark matter annihilation, and tests of substructure in gravitational lenses. The inferences derived from observations depend strongly on the assumed clustering of dark matter on the smallest scales. Unfortunately, this remains unclear from theory; while there has been progress in understanding the average density profile of isolated haloes, it remains unclear how this profile evolves once these become subhaloes. Idealized cases of single mergers have been simulated at increasing resolution, but remain susceptible to numerical effects which can cause increased mass loss [e.g. 19, 20] and decreased central densities [e.g. 23, 25]. To attempt to overcome numerical limitations, we have used a theoretically motivated, universal mass-loss model to study subhalo evolution. We quantify the effect of underestimating the central density of tidally stripped haloes on annihilation signals and lensing profiles of individual subhaloes.
For lensing mass experiments, we conclude that the details of tidal stripping are less important than assumptions about the original density profile. Thus, recent estimates of the uncertainties in mass introduced by different profiles (approximately 10 per cent) are likely reliable. On the other hand, annihilation constraints are extremely sensitive to uncertainties both in subhalo profiles and in the details of tidal stripping. In particular, our Energy Truncation model predicts annihilation signals that are an order of magnitude larger than some previous estimates, and our initial calculations indicate that the total Milky Way annihilation signal may differ by a factor of five or more depending on the subhalo model used. Future work needs to be done to fully quantify the uncertainties in current dark matter annihilation constraints in light of updated subhalo evolution models.
@@ Acknowledgments
NED acknowledges support from NSERC Canada, through a postgraduate scholarship. JET acknowledges financial support from NSERC Canada, through a Discovery Grant. Software: numpy [92], matplotlib [93], scipy [94], Colossus [95].